# AnyAnomaly: Zero-Shot Customizable Video Anomaly Detection with LVLM

### v1

@@ 4.6 Comparison with SOTA
To assess the effectiveness of AnyAnomaly in handling multiple text inputs, we conducted experiments on the VAD benchmark datasets. For performance evaluation, each anomaly class in the dataset was treated as $X$, and the maximum anomaly score among all computed scores was assigned to the corresponding segment. Table 6 presents a performance comparison with frame-centric VAD methods. Despite not being trained on VAD datasets, AnyAnomaly demonstrated a performance comparable to that of SOTA methods. Notably, it achieved a new SOTA performance of 74.5% on the UB dataset and the second-best performance of 79.7% on the ShT dataset. These results highlight the effectiveness of the proposed model, which leverages the LVLM and context-aware VQA, achieving performance on par with that of fully supervised methods.
@@ 4.7 Generalization Performance Comparison
Table 7 presents a comparison of the generalization performance of AnyAnomaly. Although STEAL-Net [5] and Jigsaw [29] achieved high accuracy in same-domain testing, their performance was significantly degraded in cross-domain settings. Specifically, on the Ave dataset, the performances of STEAL-Net and Jigsaw decreased as 87.1% → 54.3% and 92.2% → 62.9%, respectively. Similarly, on the ShT dataset, their performance decreased as 73.7% → 51.7% and 84.3% → 59.3%, respectively. This suggests that the existing OCC-based VAD models tend to overfit the training data, making them less effective when applied to new environments. For instance, ‘Too close’ where an object is in close proximity to the camera is considered anomalous in the Ave dataset but normal in the ShT dataset. Consequently, OCC-based models trained on ShT struggle to detect such anomalies.
The zero- and few-shot VAD models designed for xVAD exhibited better generalization performance than the OCC-based models. However, few-shot models depend heavily on the number of K-shot samples, whereas zero-shot models require auxiliary datasets. Training-free methods such as ZS-CLIP and ZS-ImageBind leverage strong image understanding capabilities and outperform some VAD models. Nevertheless, their performance remains limited by domain gaps. In contrast, AnyAnomaly effectively overcomes the domain gap by incorporating contextual information, achieving superior performance.
@@ A.1 Dataset Details
VAD Dataset. We used the CUHK Avenue (Ave) [16], ShanghaiTech Campus (ShT) [18], and UBnormal (UB) [1] datasets. Ave comprises of videos captured by a single camera on a university campus, containing five types of abnormal events; throwing paper, running, dancing, approaching the camera (Too close) and bicycle. ShT is a campus CCTV dataset that includes 13 different background scenes and 11 types of abnormal events; such as bicycles, cars, fighting, and jumping. UB is a synthetic dataset generated using the Cinema4D software, encompassing 29 diverse background scenes, including indoor environments, sidewalks, and etc. It provides a total of 22 abnormal events, including not only challenging-to-detect events such as smoking and stealing but also complex scenarios such as driving outside the lane and people-car accidents.
@@ A.2 Implementation Details
In a key experiment using the C-VAD datasets, we employed an efficient Chat-UniVi [13] 7B model, considering the balance between performance and speed. For the VAD dataset experiment, we utilized the effective MiniCPM-V [34] 8B model to achieve optimal performance and compared it with state-of-the-art (SOTA) models. The CLIP model used for key frames selection and context generation was ViT-B/32. For context generation, we adopted large, middle, and small window sizes of (120,120), (80,80), and (48,48), respectively. For C-Ave and Ave, the large window size was set to (240,240). All the experiments were conducted on a single NVIDIA GeForce RTX 3090 GPU.
@@ A.3 Prompt Details
Figure S1 shows the detailed prompts used in the experiments. First, a reasoning prompt is designed to obtain the chain-of-thought [30] effect by requiring a simple reason along with the anomaly score. This helps to break down the problem step-by-step, guiding the model to resolve complex issues more systematically. For example, the question “Does the image include jumping? can be divided into two steps: 1. “Is there an object related to jumping (e.g., a person)?” and 2. “Is the object performing a jumping action?” This allows object-level image analysis, leading to more refined predictions. The consideration prompt encourages the assignment of a high score even when $X$ is not central within the image. This prompt was introduced to address the issue where low scores are assigned simply because $X$ exists but is not the central element. The effectiveness of this prompt tuning is compared and analyzed in Table S1.
@@
Appendix B Additional Quantitative Evaluation
B.1 Hyperparameter Tuning
We tuned the three hyperparameters $\gamma_{1},\gamma_{2},and\gamma_{3}$ used for the final anomaly score calculation for each VAD dataset. Each hyperparameter controls the influence of the anomaly score derived from the frame, position, and temporal contexts. As shown in Table S2, the optimal hyperparameter values vary across datasets owing to differences in object sizes and abnormal events. Additionally, comparing w/o context, which does not utilize context information, and w/o tuning, where all hyperparameters were set to the same value, we observed performance improvements of 3.0%, 2.2%, and 0.7%, even without hyperparameter tuning. In contrast, the performance differences owing to hyperparameter tuning were 2.9%, 0.3%, and 0.7%, respectively. This demonstrates the effectiveness of our proposed approach in utilizing context information in VAD and proves that it achieves a strong generalization performance even without hyperparameter tuning.
B.2 Diverse LVLM Comparison
Table S3 presents the results for C-ShT and C-Ave when using various LVLMs. We evaluated the performances of four SOTA LVLMs: Chat-UniVi [13], MiniGPT-4 [38], MiniCPM-V [34], and LLAVA++ [23]. All experiments were conducted using the default settings, and ‘Pre-trained’ refers to the names of the pre-trained model weights. The experimental results demonstrate that incorporating the proposed context-aware VQA improves the performance of all LVLMs. Specifically, the use context-aware VQA leads to improvements ranging from 2.6% to 24.8%. Notably, even MiniCPM, which achieved the best performance without context-aware VQA, and showed additional improvements of 2.7% and 5.4% for C-ShT and C-Ave, respectively, when context-aware VQA was applied. This confirms that leveraging the proposed context-aware VQA is effective for C-VAD. Additionally, we observed that Chat-UniVi, with an FPS of 6.67, was the most efficient model, whereas MiniCPM-V achieved the highest performance on both datasets, scoring 90.1% and 91.0%, respectively. Therefore, as mentioned in Section A.1, Chat-UniVi was used for the C-VAD experiments and MiniCPM-V was used for the VAD dataset experiments.
Appendix C Additional Qualitative Evaluation
C.1 Context Complementarity
C.2 Anomaly Detection in Diverse scenarios
C.3 Anomaly Detection in Complex scenarios
Additionally, in Figure 4(d), we observe that the model can detect certain frames of “walking drunk” even without utilizing context information. This suggests that the strong visual reasoning capabilities of the LVLM enable VAD in complex scenarios. However, as shown in Figure 4(a)4(c), relying solely on individual frames is insufficient for fully leveraging these reasoning abilities. Therefore, the proposed context-aware VQA approach is essential for effective VAD.
Appendix D Limitations
Efficiency is crucial in VAD; therefore, we utilized the most lightweight model among the SOTA LVLMs and adopted a segment-level approach to significantly reduce the latency. However, our method still requires three inputs per segment (key frame, position context, and temporal context) and involves a reasoning process, which makes real-time analysis more challenging. Furthermore, when multiple abnormal events occur simultaneously, each event must be processed independently, which leads to a substantial increase in latency. Hence, our future studies will aim to enhance to the efficiency of the C-VAD in handling multiple abnormal events simultaneously.
Table S3: Comparison of diverse LVLMs. The model highlighted in blue represents the most efficient model for the C-VAD task, while the one highlighted in purple indicates the most effective model.
@@ Appendix D Limitations
Chat-UniVi[13] Chat-UniVi-7B 77.5 85.7 78.3 90.3 6.67
MiniGPT-4[38] LLaMA-2 Chat 7B 54.0 67.4 53.9 55.3 1.26
MiniCPM-V[34] MiniCPM-Llama3-V-2_5 (8B) 87.7 90.1 86.3 91.0 1.36
LLAVA++[23] LLaVA-Meta-Llama-3-8B-Instruct-FT 73.3 82.8 59.0 69.4 7.25

### v4

@@ 4.6 Comparison with SOTA
To assess the effectiveness of AnyAnomaly in handling multiple text inputs, we conducted experiments on the VAD benchmark datasets. For performance evaluation, each anomaly class in the dataset was treated as $X$, and the maximum anomaly score among all computed scores was assigned to the corresponding segment. Table 5 presents a performance comparison with frame-centric VAD methods. Despite not being trained on VAD datasets, AnyAnomaly demonstrated a performance comparable to that of SOTA methods. Notably, it achieved the best AUC results on the UB and UCF datasets, which cover diverse scenes and abnormal events, thereby demonstrating the effectiveness of the proposed model across various environments. Furthermore, while LLM-based methods (e.g., AnomalyRuler) require rule generation and aggregation using a few normal samples, the proposed method achieves competitive performance solely through zero-shot inference, highlighting its practical applicability.
@@ 4.7 Generalization Performance Comparison
Table 6 presents a comparison of the generalization performance of AnyAnomaly. Although STEAL-Net [astrid2021synthetic] and Jigsaw [wang2022video] achieved high accuracy in same-domain testing, their performance was significantly degraded in cross-domain settings. Specifically, on the Ave dataset, the performances of STEAL-Net and Jigsaw decreased as 87.1% → 54.3% and 92.2% → 62.9%, respectively. Similarly, on the ShT dataset, their performance decreased as 73.7% → 51.7% and 84.3% → 59.3%, respectively. This suggests that the existing OCC-based VAD models tend to overfit the training data, making them less effective when applied to new environments. For instance, ‘Too close’ where an object is in close proximity to the camera is considered anomalous in the Ave dataset but normal in the ShT dataset. Consequently, OCC-based models trained on ShT struggle to detect such anomalies.
The zero- and few-shot VAD models designed for xVAD exhibited better generalization performance than the OCC-based models. However, few-shot models depend heavily on the number of K-shot samples, whereas zero-shot models require auxiliary datasets. Training-free methods using VLMs, such as ZS-CLIP and Video-ChatGPT, leverage strong image and video understanding capabilities, outperforming some VAD models. Nevertheless, their performance is limited by domain gaps. In contrast, AnyAnomaly effectively overcomes these gaps by incorporating contextual information, achieving superior performance.
@@ 5 Conclusion
Acknowledgements
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT).(No. RS-2025-02312833)
[…]
@@ A.1 Dataset Details
VAD Dataset. We used the CUHK Avenue (Ave) [lu2013abnormal], ShanghaiTech Campus (ShT) [luo2017revisit], UBnormal (UB) [acsintoae2022ubnormal], and UCF-Crime (UCF) [sultani2018real] datasets. Ave comprises of videos captured by a single camera on a university campus, containing five types of abnormal events; throwing paper, running, dancing, approaching the camera (Too close) and bicycle. ShT is a campus CCTV dataset that includes 13 different background scenes and 11 types of abnormal events; such as bicycles, cars, fighting, and jumping. UB is a synthetic dataset generated using the Cinema4D software, encompassing 29 diverse background scenes, including indoor environments, sidewalks, and etc. It provides a total of 22 abnormal events, including not only challenging-to-detect events such as smoking and stealing but also complex scenarios such as driving outside the lane and people-car accidents. UCF is a large-scale dataset containing 1,900 untrimmed real-world surveillance videos, covering 13 anomaly categories such as assault, burglary, explosion, and stealing.
@@ A.2 Implementation Details
In a key experiment using the C-VAD datasets, we employed an efficient Chat-UniVi [jin2024chat] 7B model, considering the balance between performance and speed. For the VAD dataset experiment, we utilized the effective MiniCPM-V [yao2024minicpm] 8B model to achieve optimal performance and compared it with state-of-the-art (SOTA) models. At the time, both models represented a well-validated choice for context-aware VQA. Additional results with Qwen2.5-VL [Qwen2.5-VL] are reported in Table S4 for further comparison. The CLIP model used for key frames selection and context generation was ViT-B/32. For context generation, we adopted large, middle, and small window sizes of (120,120), (80,80), and (48,48), respectively. For C-Ave and Ave, the large window size was set to (240,240). All the experiments were conducted on a single NVIDIA GeForce RTX 3090 GPU.
@@ A.3 Prompt Details
Figure S1 shows the detailed prompts used in the experiments. First, a reasoning prompt is designed to obtain the chain-of-thought [wei2022chain] effect by requiring a simple reason along with the anomaly score. This helps to break down the problem step-by-step, guiding the model to resolve complex issues more systematically. For example, the question “Does the image include jumping? can be divided into two steps: 1. “Is there an object related to jumping (e.g., a person)?” and 2. “Is the object performing a jumping action?” This allows object-level image analysis, leading to more refined predictions. The consideration prompt encourages the assignment of a high score even when $X$ is not central within the image. This prompt was introduced to address the issue where low scores are assigned simply because $X$ exists but is not the central element. The effectiveness of this prompt tuning is compared and analyzed in Table S1.
@@
Table S2: Comparison on segment length
Segment length C-ShT C-Ave FPS
Baseline 78.01 79.43 0.96
8 83.83 83.96 2.67
16 83.45 87.45 4.49
24 85.72 90.27 6.67
32 82.50 85.94 8.45
@@
Table S4: Comparison of diverse LVLMs. The model highlighted in blue represents the most efficient model for the C-VAD task, while the one highlighted in purple indicates the most effective model. For further comparison, additional experiments were conducted using Qwen-based models. *: Experiment conducted using vLLM.
@@
Chat-UniVi[jin2024chat] Chat-UniVi-7B 77.5 85.7 78.3 90.3 6.67
MiniGPT-4[zhu2023minigpt] LLaMA-2 Chat 7B 54.0 67.4 53.9 55.3 1.26
MiniCPM-V[yao2024minicpm] MiniCPM-Llama3-V-2_5 (8B) 87.7 90.1 86.3 91.0 1.36
LLAVA++[hanoona2024LLaVA++] LLaVA-Meta-Llama-3-8B-Instruct-FT 73.3 82.8 59.0 69.4 7.25
Qwen2.5-VL[Qwen2.5-VL] Qwen2.5-VL-3B-Instruct 89.0 90.2 78.0 87.0 11.18
Qwen2.5-VL*[Qwen2.5-VL] Qwen2.5-VL-3B-Instruct 88.6 90.2 78.3 88.1 34.78
Qwen2.5-VL*[Qwen2.5-VL] Qwen2.5-VL-7B-Instruct 93.0 95.5 86.9 92.4 24.08
Appendix B Additional Quantitative Evaluation
B.1 Segment length and FPS
Table S2 presents the performance comparison and FPS based on different segment lengths. The baseline segment length was set to 1. It was observed that deriving anomaly scores at the segment level yields superior performance compared to the baseline, which relies on a single frame. The highest AUC performance was achieved when the segment length is set to 24, reaching 85.72% and 90.27% for C-ShT and C-Ave, respectively. However, excessively long segment length introduces irrelevant information into the temporal context, leading to a decrease in accuracy. Furthermore, performing VAD at the segment level resulted in a 594% improvement in the FPS compared with the baseline.
B.2 Hyperparameter Tuning
We tuned the three hyperparameters $\gamma_{1},\gamma_{2}$, and $\gamma_{3}$ used for the final anomaly score calculation for each VAD dataset. Each hyperparameter controls the influence of the anomaly score derived from the frame, position, and temporal contexts. As shown in Table S3, the optimal hyperparameter values vary across datasets owing to differences in object sizes and abnormal events. Additionally, comparing w/o context, which does not utilize context information, and w/o tuning, where all hyperparameters were set to the same value, we observed performance improvements of 3.0%, 2.2%, 0.7%, and 2.7%, even without hyperparameter tuning. In contrast, the performance differences owing to hyperparameter tuning were 2.9%, 0.3%, 0.7% and 0.2%, respectively. This demonstrates the effectiveness of our proposed approach in utilizing context information in VAD and proves that it achieves a strong generalization performance even without hyperparameter tuning.
B.3 Diverse LVLM Comparison
Table S4 presents the results for C-ShT and C-Ave when using various LVLMs. We evaluated the performances of four SOTA LVLMs: Chat-UniVi [jin2024chat], MiniGPT-4 [zhu2023minigpt], MiniCPM-V [yao2024minicpm], and LLAVA++ [hanoona2024LLaVA++]. All experiments were conducted using the default settings, and ‘Pre-trained’ refers to the names of the pre-trained model weights. The experimental results demonstrate that incorporating the proposed context-aware VQA improves the performance of all LVLMs. Specifically, the use context-aware VQA leads to improvements ranging from 2.6% to 24.8%. Notably, even MiniCPM, which achieved the best performance without context-aware VQA, and showed additional improvements of 2.7% and 5.4% for C-ShT and C-Ave, respectively, when context-aware VQA was applied. This confirms that leveraging the proposed context-aware VQA is effective for C-VAD. Additionally, we observed that Chat-UniVi, with an FPS of 6.67, was the most efficient model, whereas MiniCPM-V achieved the highest performance on both datasets, scoring 90.1% and 91.0%, respectively. Therefore, as mentioned in Section A.1, Chat-UniVi was used for the C-VAD experiments and MiniCPM-V was used for the VAD dataset experiments.
B.4 Additional Experiments with vLLM
With recent advances in LLM inference libraries such as vLLM [kwon2023efficient], latency issues can be largely mitigated. vLLM supports continuous batching and KV-caching, which substantially improve inference speed. Our method is not restricted to a specific LVLM and can be applied to various models supported by vLLM.
To evaluate efficiency, we conducted experiments by batching the key frame, $PC$, and $TC$ during C-VAD. The experiments were conducted using Qwen2.5-VL [Qwen2.5-VL], one of the representative models supported by vLLM. The results, presented in the lower part of Table S4, show that using vLLM maintained the AUC while achieving a 211% improvement in FPS (from 11.18 to 34.78). Furthermore, even with a 7B model, higher FPS was observed compared to other models. These findings suggest that AnyAnomaly has strong potential for real-time applications.
Appendix C Additional Qualitative Evaluation
C.1 Further Analysis on Contexts
To provide a deeper understanding of the role of $PC$ and $TC$, we present both qualitative analyses of their effectiveness and representative failure cases.
Figure S2 illustrates the improvements achieved by incorporating $PC$ and $TC$. Without $PC$, the motorcycle appeared too small to be detected, resulting in an anomaly score of 0. With $PC$ applied, however, the relevant region was emphasized, and the score increased to 0.8 despite the object’s small size (top row). Likewise, without $TC$, the model misinterpreted skateboarding as walking and assigned a low anomaly score of 0.1. By integrating temporal information, the model captured motion cues such as positional changes and the enlarged appearance of the skateboard, correctly identifying skateboarding with a score of 0.75 (bottom row). These results indicate that our context-aware VQA is more effective than conventional VQA.
Nevertheless, Figure S3 demonstrates that $PC$ and $TC$ can also introduce errors under certain conditions. In the top row, before applying $PC$, the LVLM detected the abnormal action of throwing a bag with an anomaly score of 0.7. After applying $PC$, however, the model attended more to the person than the bag, making the action ambiguous and reducing the score to 0.3. In the bottom row, without $TC$, a normal walking scene yielded a score of 0, but after applying TC, walking was misinterpreted as a hard negative, raising the score to 0.4. While these failure cases exist, they provide valuable insights into the challenges of context-aware VQA and guide future directions for improvement.
Figure S2: Anomaly score comparison with context visualization (success cases)
Figure S3: Anomaly score comparison with context visualization (failure cases)
@@ C.1 Further Analysis on Contexts
C.2 Context Complementarity
C.3 Anomaly Detection in Diverse scenarios
C.4 Anomaly Detection in Complex scenarios
Additionally, in Figure 6(d), we observe that the model can detect certain frames of “walking drunk” even without utilizing context information. This suggests that the strong visual reasoning capabilities of the LVLM enable VAD in complex scenarios. However, as shown in Figure 6(a)6(c), relying solely on individual frames is insufficient for fully leveraging these reasoning abilities. Therefore, the proposed context-aware VQA approach is essential for effective VAD.
Appendix D Discussion
D.1 Comparison with traditional VAD
Traditional VAD methods and our zero-shot C-VAD each have distinct strengths and limitations. Traditional VAD detects anomalies as deviations from learned normal patterns, requiring no prior knowledge of specific anomaly types and delivering strong performance within the trained environment. However, it often exhibits poor generalization to unseen environments and typically necessitates retraining. In contrast, C-VAD requires prior knowledge of anomaly types but removes the need for retraining or additional data collection even when the definition of “normal” varies across users or environments. This makes it a practical and cost-effective solution for real-world applications. We anticipate that, with continued advances in LVLM technology, the proposed C-VAD will become even more effective in the future.
D.2 Limitation
Efficiency is crucial in VAD; therefore, we adopted the most lightweight model among the SOTA LVLMs and employed a segment-level approach to reduce latency. However, when multiple abnormal events occur simultaneously, each event must be processed independently, which increases latency. In future work, we plan to improve the efficiency of C-VAD in handling multiple abnormal events concurrently.